# The Promise of Time-Series Foundation Models for Agricultural Forecasting: Evidence from Marketing Year Average Prices

### v1

@@ 2.3 USDA Official MYA Price Forecast
where $\text{Futures}_{t}$ is the nearby futures contract settlement price for month $t$, and $\overline{\text{Basis}}_{t}$ is the historical average basis (cash-futures differential) for month $t$, typically computed as a 5-year moving average for corn, soybeans, and wheat, or a 7-year Olympic average for cotton. The basis reflects local supply-demand conditions, transportation costs, and storage premiums that cause cash prices to deviate from futures prices. By combining liquid futures markets (which aggregate information about expected supply and demand) with historical basis patterns (which capture local market structure), this methodology leverages both forward-looking market expectations and historical regularities. [^3] This creates information asymmetry in our comparison: USDA forecasts incorporate forward-looking futures prices, while our models use only historical cash prices. We discuss the implications of this design choice in Section 5. Briefly, this represents a more stringent test for our models – outperforming futures-based forecasts using only backward-looking information demonstrates genuine forecasting ability rather than simply incorporating the same market signals.
Monthly forecasts are then aggregated to MYA using marketing percentage weights as in Equation (1). Months with actual prices use observed values; remaining months use futures-plus-basis forecasts. This offical MYA forecast serves as our primary benchmark as it is institutional agricultural price forecasting used by USDA World Agricultural Outlook Board (WAOB) in World Agricultural Supply and Demand Estimates (WASDE) reports.
To ensure fair comparison with USDA, we use their final forecast made before the marketing year begins (approximately 4 days prior to the start date). Our models forecast 12 monthly prices using only information available at that same point in time, then aggregate to MYA price using the same marketing percentage weights. This ensures our models operate under a restrictive information set, with no future information leakage.
@@ 3.1.1 Traditional Time Series Models
where $B$ is the backshift operator ($BP_{t}=P_{t-1}$), $s=12$ is the seasonal period, $\phi(B)$ and $\Phi(B^{s})$ are autoregressive polynomials for non-seasonal and seasonal components, $\theta(B)$ and $\Theta(B^{s})$ are moving average polynomials, and the differencing operators $(1-B)^{d}$ and $(1-B^{s})^{D}$ remove non-seasonal and seasonal trends respectively. For $h$-step ahead forecasting, SARIMA generates predictions recursively: the model first forecasts $\hat{P}_{t+1}$, then uses this forecast as input to predict $\hat{P}_{t+2}$, continuing until reaching horizon $h$. The algorithm selects optimal orders based on information criterion separately for each commodity-split combination, balancing model fit against complexity. SARIMA benefits from a well-established statistical foundation with interpretable parameters and the ability to handle non-stationarity through differencing, but it assumes linear relationships, which is sensitive to outliers, and requires sufficient historical data for reliable parameter estimation.
@@ 3.4 Cross-Validation Design and Hyperparameter Selection
With 4 commodities and 16 splits, we generate 64 split-commodity combinations for model comparison and 1,088 total forecasts across all models ($17\times 64$). For MYA price prediction comparison, data availability varies by commodity, yielding 49 USDA-comparable forecasts. See Appendix C for complete cross-validation details including split definitions, hyperparameter specifications, and optimization experiments.
@@ 4 Empirical Results
We evaluate model performance in two applications: monthly price forecasting across all 17 models, and MYA forecasts benchmarked against official USDA projections. This section presents results for both sets of evaluations, identifying top performers and analyzing patterns across commodities.
@@ 4.1.1 Overall Model Rankings
Notes: Monthly forecast performance averaged across 768 forecasts ($64$ splits $\times$ 1212 months, with some variation by commodity).
@@ 4.2 Evaluation 2: MYA Forecasting vs USDA Benchmark
We now evaluate MYA forecasts against USDA operational benchmarks using 49 matched marketing years where USDA forecasts are available (corn and soybeans: 14 years from 2009; wheat: 15 years from 2003; cotton: 6 years from 2019). Models aggregate their 12-month-ahead forecasts to MYA using marketing percentage weights, enabling direct comparison with USDA’s forecasts. For each model, we compute commodity-specific MAE by averaging absolute errors across all matched years.
@@ 4.2.1 Best Model Performance by Commodity
Table 5 presents the best-performing model for each commodity against USDA operational forecasts. The best models achieve substantial improvements over USDA: 52.9% on corn (Time-MoE), 55.2% on soybeans (Time-MoE), 25.5% on wheat (Time-MoE), and 5.1% on cotton (Seasonal Naive). Time-MoE dominates on major row crops, capturing complex price dynamics that futures-based forecasts miss. Cotton shows the smallest gain, where Seasonal Naive performs best though Time-MoE achieves nearly identical accuracy (MAE 0.103 vs 0.102). Overall, Time-MoE improves 45.4% over USDA across all commodities, demonstrating consistent strong performance.
Table 5: Best Model Performance vs USDA by Commodity
@@
Corn (14 years) 0.995 Time-MoE 0.469 +52.9%
Soybeans (14 years) 1.528 Time-MoE 0.685 +55.2%
Wheat (15 years) 0.911 Time-MoE 0.679 +25.5%
Cotton (6 years) 0.108 Seasonal Naive 0.102 +5.1%
Overall 0.886 Time-MoE 0.484 +45.4%
Notes: Best model selected based on lowest MAE for each commodity. USDA forecasts made approximately 4 days before marketing year starts. Improvement = (USDA MAE - Model MAE) / USDA MAE.
@@ 4.2.2 Comprehensive Model Comparison Across Commodities
Table 6 presents MYA forecast performance for all 17 models across four commodities. Foundation models dominate, with Time-MoE achieving best performance on corn, soybeans, and wheat. Traditional models show competitive performance on some commodities, with Naive and SARIMA achieving significant improvements over USDA on corn and soybeans. Deep learning models consistently underperform across all commodities despite extensive hyperparameter optimization, ranking in the bottom tier. Machine learning models show mixed results: XGBoost ranks second on corn but struggles on soybeans and wheat.
Diebold-Mariano tests (reported as significance stars) again confirm that foundation models’ advantages over USDA are statistically robust for corn and soybeans, where larger sample sizes (n=14 each) provide sufficient statistical power. Wheat shows no significant results despite Time-MoE achieving 25.5% improvement: the HAC standard error of the test statistic is relatively large, yielding a p-value of 0.14. This reflects high variance in wheat price forecast errors rather than lack of genuine performance differences. Cotton shows no significant results because USDA’s futures-based forecasts are already highly accurate (MAE \$0.108), leaving little room for statistically significant improvements. Time-MoE achieves statistical significance on both major row crops (corn and soybeans), demonstrating that its performance gains are not due to chance.
Table 6: Model Performance by Commodity - All Models (MAE, \$/unit)
@@
Time-MoE 0.469∗∗∗ 0.685∗∗∗ 0.679 0.103
TimesFM 2.5 0.589∗∗∗ 0.862 0.798 0.135
Chronos 0.620∗∗ 0.749∗∗∗ 0.826 0.159
Chronos-2 0.633∗∗∗ 0.738∗∗∗ 0.846 0.126
Moirai-2 0.571∗∗∗ 0.864∗∗ 0.859 0.133
@@
Random Forest 0.576∗∗∗ 0.918 0.907 0.139
XGBoost 0.538∗∗∗ 1.026 0.874 0.122
LSTM 0.717 1.138 0.771 0.118
@@
Naive 0.646∗∗∗ 0.981 0.854 0.144
Seasonal Naive 0.813∗∗ 1.406 1.129 0.102
SARIMA 0.813 0.890∗∗ 0.947 0.144
Exp Smoothing 0.856 1.153 0.866 0.141
STL 1.036 1.532 0.892 0.133
Prophet 1.296 2.124 1.197 0.119
Deep Learning (trained from scratch)
N-BEATS 0.760 1.309 0.910 0.144
TFT 0.916 1.202 1.014 0.105
DeepAR 1.123 1.892 1.305 0.150
Notes: Bold values indicate best performance for each commodity. Significance stars indicate Diebold-Mariano test results comparing each model to USDA ERS baseline: p<<0.10, ∗∗ p<<0.05, ∗∗∗ p<<0.01. Stars appear only when the model outperforms USDA. DM tests use absolute error loss differentials with Newey-West HAC standard errors. Although each MYA forecast aggregates 12 monthly predictions, the annual MYA errors are non-overlapping across marketing years, so we set $h=1$ for lag selection.
@@ 5.1 Practical Implications for Agricultural Policy
Marketing Year Average price forecasts play a central role in administering Farm Bill safety net programs, where Price Loss Coverage (PLC) and Agriculture Risk Coverage (ARC) payments depend on MYA price realizations (Zulauf and Schnitkey, 2014; Schnitkey et al., 2019). Time-MoE achieves 45.4% improvement over USDA’s operational forecasts, with particularly large gains on corn (52.9%) and soybeans (55.2%) – the two largest U.S. field crops representing over 175 million acres annually. A 50% reduction in forecast error translates to more reliable program payment estimates, potentially affecting billions of dollars in farm safety net expenditures.
Operational deployment by USDA would require addressing explainability concerns, as foundation models operate as black boxes. A hybrid approach could use foundation models to generate primary forecasts while maintaining traditional methods (SARIMA, futures-basis) to provide interpretable validation checks and explain forecast drivers to policy stakeholders.
@@ 5.2 Why Deep Learning Models Fail on Limited Agricultural Data
5.3 Comparison with USDA’s Information Set
Our models outperform USDA despite using only historical cash prices, while USDA incorporates forward-looking futures prices reflecting traders’ expectations. This restricted information set establishes a lower bound on foundation model performance. The substantial improvements (45.4% overall) suggest two possible explanations: foundation models capture price dynamics not fully reflected in futures markets, or the historical basis adjustments in USDA’s futures-plus-basis methodology introduce systematic bias.
Given that futures markets efficiently aggregate information, the latter explanation may be more plausible. USDA’s methodology uses 5-7 year average basis (cash-futures differential) to adjust futures prices, assuming basis patterns remain stable. However, basis reflects local market conditions – storage costs, transportation infrastructure, regional supply-demand imbalances – that can shift due to infrastructure changes, policy reforms, market consolidation, or technological improvements in logistics. Historical basis may therefore fail to predict future basis realizations, particularly during periods of structural change or market disruption.

### v2

@@ 2.2 USDA ERS Season-Average Price Forecasts
where $\widehat{P}_{k,t}$ is the predicted farm price in calendar month $k$ and year $t$, $F_{h(k),t,i}$ is the futures price for the nearby contract expiring in month $h(k)$ observed at forecast horizon $i$, and $\bar{B}_{k,t}$ is the historical average basis. [^4] For corn, the nearby contract mapping follows the harvest cycle: December futures for October-November prices, March futures for December-February, May futures for March-April, July futures for May-June, and September futures for July-August. Similar mappings apply to other commodities based on their respective marketing years and contract availability. This methodology forecasts each month directly using the corresponding nearby futures contract plus historical basis, avoiding error accumulation across the 12-month horizon. The basis reflects local supply-demand conditions, transportation costs, and storage premiums that cause cash prices to deviate from futures prices. The basis averages use 5-year moving windows for corn, soybeans, and wheat, and 7-year Olympic averages for cotton to address structural changes while maintaining forecast reliability. [^5] This creates information asymmetry in our comparison: USDA forecasts incorporate forward-looking futures prices, while our models use only historical cash prices. We discuss the implications of this design choice in Section 5. Briefly, this represents a more stringent test for our models – outperforming futures-based forecasts using only backward-looking information demonstrates genuine forecasting ability rather than simply incorporating the same market signals.
Monthly forecasts are then aggregated to the Marketing Year Average (MYA), which represents the weighted average price over the 12-month marketing year:
$\widehat{\text{MYA}}_{t}$ $=\sum_{k=1}^{12}\widehat{P}_{k,t}\times w_{k},$ (2)
where $\widehat{P}_{k,t}$ is the monthly price predicted by Equation (1) and $w_{k}$ is the monthly marketing percentage – the proportion of annual production marketed in month $k$, with $\sum_{k=1}^{12}w_{k}=1$. The actual MYA is calculated using actual farmer-received prices and actual marketing percentages within that marketing year. Marketing percentages reflect the typical temporal pattern of commodity sales throughout the year. For example, corn marketing percentages are highest immediately post-harvest (September–November: 45% of annual sales) and decline through the marketing year as on-farm storage is depleted. For forecasting purposes, each month’s marketing percentage uses the 5-year average of that same month’s historical percentage for corn, soybeans, and wheat, and a 7-year Olympic average (dropping the highest and lowest values) for cotton.
@@ 3.1.1 Traditional Time Series Models
where $B$ is the backshift operator ($BP_{t}=P_{t-1}$), $s=12$ is the seasonal period, $\phi(B)$ and $\Phi(B^{s})$ are autoregressive polynomials for non-seasonal and seasonal components, $\theta(B)$ and $\Theta(B^{s})$ are moving average polynomials, and the differencing operators $(1-B)^{d}$ and $(1-B^{s})^{D}$ remove non-seasonal and seasonal trends respectively. For $h$-step ahead forecasting, SARIMA generates predictions recursively: the model first forecasts $\hat{P}_{t+1}$, then uses this forecast as input to predict $\hat{P}_{t+2}$, continuing until reaching horizon $h$. The algorithm selects optimal orders based on information criterion separately for each commodity-split combination, balancing model fit against complexity. SARIMA benefits from a well-established statistical foundation with interpretable parameters and the ability to handle non-stationarity through differencing, but it assumes linear relationships, is sensitive to outliers, and requires sufficient historical data for reliable parameter estimation.
@@ 3.4 Cross-Validation Design and Hyperparameter Selection
With 4 commodities and 16 splits, we generate 64 split-commodity combinations for model comparison and 1,088 total forecasts across all models ($17\times 64$). For MYA price prediction comparison, data availability varies by commodity, yielding 49 USDA-comparable forecasts. To ensure fair comparison with USDA ERS, we use their second-to-last weekly-updated forecast before the marketing year begins to avoid any overlap with the marketing year. Our models forecast 12 monthly prices using only information available before the marketing year starts, then aggregate to SAP using the same marketing percentage weights. This ensures our models operate under a restrictive information set, with no future information leakage. See Appendix C for complete cross-validation details including split definitions, hyperparameter specifications, and optimization experiments.
@@ 4 Empirical Results
We evaluate model performance in two applications: monthly price forecasting across all 17 models, and MYA forecasts benchmarked against USDA ERS SAP prediction. This section presents results for both sets of evaluations, identifying top performers and analyzing patterns across commodities.
@@ 4.1.1 Overall Model Rankings
Notes: Monthly forecast performance averaged across 768 forecasts ($64$ splits $\times$ $12$ months, with some variation by commodity).
@@ 4.1.3 Statistical Significance of Performance Differences
To verify that foundation models learn genuine forecasting patterns rather than memorizing historical sequences, we evaluate all 17 models on 400 synthetically generated price series matching the statistical properties of agricultural data (Appendix D). Foundation models maintain top performance on synthetic data, with Chronos-2, Moirai-2, and TimesFM ranking 1-3, confirming they have learned generalizable time series patterns rather than memorizing specific historical sequences. This validation addresses concerns about potential data contamination in foundation model pre-training corpora.
@@ 4.2 Evaluation 2: MYA Forecasting vs USDA Benchmark
We evaluate MYA forecasts against USDA ERS forecast, which is an important input to the prediction in WASDE report. ERS forecast represent a rather competitive benchmark to compare with for two fundamental reasons. First, USDA’s futures-plus-basis methodology incorporates forward-looking market information unavailable to our models. Futures prices aggregate expectations from thousands of market participants regarding supply conditions, demand shifts, weather patterns, trade policy developments, and other fundamental factors (Fama, 1970; Tomek, 1997). This information set extends well beyond the historical price patterns our models observe. Second, USDA forecasts each month directly using the corresponding nearby futures contract (December futures for October-November prices, March futures for December-February prices, and so forth) rather than iteratively. This direct approach avoids the error accumulation inherent in sequential forecasting across the 12-month horizon. These dual advantages (superior information and methodology) establish USDA as a particularly demanding benchmark. Consequently, models that outperform USDA demonstrate an ability to extract predictive patterns from historical prices that futures markets do not fully incorporate.
Our evaluation spans marketing years (MY) 2017-2024 excluding the COVID shock year (2020), yielding 7 years of data for all commodities to ensure balanced comparison. [^13] We focus on MY 2017-2024 for two reasons. First, this period provides a balanced panel where all four commodities have complete data, enabling fair cross-commodity comparison. Cotton MYA forecasts only became available from MY 2017 onward. Second, we exclude MY 2020 (September 2020 - August 2021) because the COVID-19 pandemic created unprecedented market disruptions that distort forecast errors. Including this outlier year would mask performance patterns in normal market conditions. Our focus on MY 2017-2024 is driven by data availability constraints rather than model performance considerations. This period includes substantial market volatility (2018-2019 U.S.-China trade war, 2019 Midwest flooding, 2022-2023 Ukraine war impacts), providing a stringent test of model robustness across diverse market conditions. While longer evaluation periods would strengthen external validity, extending backward would require either dropping cotton or analyzing an unbalanced panel, complicating cross-commodity comparisons. We use USDA’s second-to-last forecast before each marketing year starts to eliminate ambiguity regarding week boundaries. Models aggregate their 12-month forecasts to MYA using 5-year averaged marketing percentage weights (7-year Olympic average for cotton), replicating USDA methodology detailed in Section 2.2. For each model, we compute commodity-specific MAE by averaging absolute errors across matched years.
@@ 4.2.1 Best Model Performance by Commodity
Table 5 presents the best-performing model for each commodity. Time-MoE achieves substantial improvements on corn (18.5%) and wheat (54.9%), while TimesFM performs best on soybeans (6.9%). Cotton remains challenging, with USDA maintaining superior accuracy.
Table 5: Best Model Performance vs USDA by Commodity (MY 2017-2024, excl 2020)
@@
Corn 0.328 Time-MoE 0.267 +18.5%
Soybeans 0.396 TimesFM 0.369 +6.9%
Wheat 1.201 Time-MoE 0.542 +54.9%
Cotton 0.045 Seasonal Naive 0.094 -110.5%
Notes: Best model selected based on lowest MAE for each commodity. Improvement = (USDA MAE - Model MAE) / USDA MAE. Negative improvement indicates USDA outperforms the model.
@@ 4.2.2 Comprehensive Model Comparison Across Commodities
Table 6 presents MYA forecast performance for all 17 models across four commodities, expressed as ratios relative to USDA ERS (values below 1.00 indicate superior performance). Foundation models demonstrate strong performance, with Time-MoE achieving the best results on corn (0.82) and wheat (0.45). TimesFM performs best on soybeans (0.93), representing the only model across all categories to outperform USDA on this commodity. Traditional and machine learning models show mixed results, with several achieving competitive performance on wheat. Deep learning models trained from scratch show strong performance on wheat, with LSTM achieving 0.55 (45% improvement), though they underperform on other commodities despite extensive hyperparameter optimization.
Table 6: Model Performance by Commodity - Relative to USDA ERS
@@
Time-MoE 0.82 1.18 0.45∗∗ 2.15
TimesFM 1.08 0.93 0.63∗∗ 2.99
Chronos 0.97 1.18 0.69 3.41
Chronos-2 1.31 1.01 0.65∗∗ 2.77
Moirai-2 1.14 1.21 0.68∗∗ 2.90
@@
Random Forest 1.20 1.48 0.61 3.09
XGBoost 1.17 1.28 0.58 2.87
Deep Learning
LSTM 1.65 1.63 0.55 2.39
N-BEATS 1.52 1.72 0.64∗∗ 3.24
TFT 2.17 2.09 0.99 2.11
DeepAR 2.23 4.23 0.94 3.44
@@
Naive 1.44 1.46 0.73 3.19
Seasonal Naive 1.95 2.99 1.05 2.11
SARIMA 1.65 1.71 0.76 2.91
Exp Smoothing 2.54 1.66 0.82 2.98
STL 2.61 2.96 1.02 2.46
Prophet 3.72 5.13 1.19 2.71
Notes: Values show model MAE relative to USDA ERS baseline (model MAE / ERS MAE). Bold values indicate model outperforms ERS (ratio << 1.00). Significance stars from Diebold-Mariano tests: p<<0.10, ∗∗ p<<0.05, ∗∗∗ p<<0.01. Stars appear only when model outperforms ERS. DM tests use Newey-West HAC standard errors with h=1 lag (annual MYA errors are non-overlapping across marketing years). USDA forecasts use second-to-last forecast before marketing year starts. All models evaluated using same test years for fair comparison.
The results reveal substantial heterogeneity across commodities. Wheat exhibits the strongest performance, with 14 out of 17 models outperforming USDA, led by Time-MoE at 0.45 (55% improvement). Corn shows modest improvements, with Time-MoE (0.82) and Chronos (0.97) beating USDA. Soybeans remains competitive, with only TimesFM (0.93) outperforming USDA. Cotton proves most challenging, with all models substantially worse than USDA (ratios 2.11-3.44), reflecting the exceptional accuracy of USDA’s futures-based forecasts for this commodity. These results are particularly noteworthy given that our models use only historical prices while USDA incorporates forward-looking futures market information. The fact that pure time series models can outperform futures-based benchmarks on major commodities demonstrates their ability to extract predictive patterns not fully reflected in market expectations.
Our evaluation focuses on pre-season forecasts generated before the marketing year begins, with matched USDA ERS forecasts. This design choice reflects practical forecasting constraints: policymakers and farmers require advance price projections to inform planting decisions and program enrollment. Research on fixed-event forecasts documents that accuracy varies substantially by forecast horizon, with errors declining as more information becomes available closer to the target period (Isengildina-Massa et al., 2024). Our single-horizon design provides policy-relevant comparisons at the most challenging forecast timing, though future work could examine whether foundation model advantages persist across multiple horizons.
@@ 5.1 Practical Implications for Agricultural Policy
WASDE SAP forecasts play a central role in administering Farm Bill safety net programs, where Price Loss Coverage (PLC) and Agriculture Risk Coverage (ARC) payments depend on MYA price realizations (Hoffman and Meyer, 2018). Time-MoE achieves strong performance relative to USDA ERS forecasts on recent data, with particularly notable results on wheat (55% improvement) and corn (18% improvement) – the two largest U.S. field crops representing over 175 million acres annually. These improvements could translate to more reliable program payment estimates, potentially affecting billions of dollars in farm safety net expenditures.
Operational deployment by USDA would require addressing explainability concerns, as foundation models operate as black boxes. A hybrid approach could use foundation models to generate primary forecasts while maintaining traditional methods (SARIMA, futures-basis) to provide interpretable validation checks and explain forecast drivers to policy stakeholders. For operational implementation, USDA could generate foundation model forecasts as complementary inputs to their existing futures-based methodology. When the two approaches diverge substantially, this signals either unusual market conditions or potential basis forecast errors, triggering additional analyst review. The modest computational requirements (sub-second inference after one-time model download) make this operationally feasible without significant infrastructure investment.
Beyond farm policy applications, our findings have broader implications for commodity market participants. Previous research demonstrates that futures-based forecasts provide viable alternatives to USDA forecasts for private sector users (Colino and Irwin, 2010; Hoffman et al., 2015; Irwin et al., 1994; Irwin and Good, 2015; Manfredo and Sanders, 2004; Poghosyan et al., 2025). Our results suggest that foundation models, requiring only historical prices and no specialized market data or futures market access, offer an accessible forecasting tool for agricultural businesses, commodity traders, and financial institutions. Companies using commodities for portfolio diversification or hedging strategies can leverage these models to generate reliable price projections without the data infrastructure and market access required for futures-based approaches.
It is worth noting that foundation model advantages vary by commodity: strong for wheat and corn, modest for soybeans, absent for cotton. This suggests commodity-specific model selection rather than universal application.
@@ 5.2 Why Deep Learning Models Fail on Limited Agricultural Data
5.3 Why Foundation Models Outperform Futures-Based Forecasts
Foundation models achieve substantial improvements over USDA ERS forecasts on wheat (55%) and corn (18%), modest gains on soybeans (7%), but underperform on cotton. This pattern is particularly striking given that our models use only historical cash prices while USDA incorporates forward-looking futures prices from actively traded markets. Under market efficiency, futures prices should aggregate all available information, making them difficult to outperform. Two explanations could account for foundation model success: either they capture price dynamics not fully reflected in futures markets, or systematic errors in USDA’s basis adjustment methodology create exploitable forecast biases.
The latter explanation appears more plausible. USDA’s futures-plus-basis methodology assumes that historical basis patterns (5-7 year averages) reliably predict future basis realizations. However, basis reflects local market conditions – storage costs, transportation infrastructure, regional supply-demand imbalances – that evolve over time. Infrastructure improvements, policy reforms, market consolidation, and logistics technology changes can alter basis patterns, making historical averages poor predictors during periods of structural change.
Empirical evidence supports this interpretation. Hoffman et al. (2015) found that WASDE corn price projections outperformed futures-adjusted forecasts in only 4 out of 16 forecasting periods, suggesting substantial room for improvement. Isengildina-Massa et al. (2024) document systematic underestimation bias in USDA price forecasts for corn, soybeans, and wheat, likely reflecting analysts’ tendency to underestimate long-term growth rates. Poghosyan et al. (2025) demonstrate that incorporating current basis deviations from historical averages significantly improves cotton price forecast accuracy, confirming that fixed historical basis averages miss important market information. Foundation models may exploit similar patterns by learning adaptive relationships directly from cash price dynamics rather than relying on fixed historical basis adjustments. This mechanism explains why foundation models excel on wheat and corn, where basis patterns have experienced more structural change, while struggling on cotton, where USDA’s 7-year Olympic average basis methodology proves more stable and accurate.